OpenStudy (anonymous):

-10(x + 1) ≤ -12x + 7

OpenStudy (anonymous):

Okay this is what i got: -10(x + 1) ≤ -12x + 7 -10x - 9 ≤ -12x + 7 -9 ≤ 2x + 7 -16 ≤ 2x x = -8 Is that what you are looking for?

OpenStudy (anonymous):

this is just like solving for x=something. get all the x on one side and everything else on the other. A few special rules are, if you multiply by a negative across the whole equation, then the inequality switches.

OpenStudy (anonymous):

x doesn't equal -8! \[\large x \ge -8\]

OpenStudy (anonymous):

oh yeah i see what u mean but i got the -8 part right lol ^.^ thanks

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

how did you get the -16

OpenStudy (anonymous):

everything i do i dont get -16

OpenStudy (anonymous):

see the -9 ≤ 2x + 7? okay, we subtract 7 from -9, which gives -9 - 7 which ends up adding instead of subtracting becuz they are both negative numbers. so just like 9 +7 = 16, but instead since they are both negative, it will be -9 - 7 = -16.

OpenStudy (anonymous):

Get it?

OpenStudy (anonymous):

I see a problem in the math from up above. it's -10x-10, not -10x-9...

OpenStudy (anonymous):

you guys are really losing me

OpenStudy (anonymous):

alright, step by step. \[\huge -10x-10 \le -12x + 7\] \[\large -10x \cancel {- 10 + 10} + 12x \le cancle{-12x} + 7 +10 \cancel{+12x}\]

OpenStudy (anonymous):

Doh typo.. but you see what happens right?

OpenStudy (anonymous):

how did the 12 come one th other side

OpenStudy (anonymous):

added it on both sides. Just like solving for x.

OpenStudy (anonymous):

s now its -10+12x ≤ -12+7+10

OpenStudy (anonymous):

\[\large -10x+12x \le 7+10\] It's this. you see where it says "cancle" ? that was my typo and it should have a line through the -12x.

OpenStudy (anonymous):

x\[x \le17/2\]

OpenStudy (anonymous):

-10(x + 1) ≤ -12x + 7 -10x -10+10+12x ≤cancel -12x +7+10+12x -10x+12x≤7+10 -10x+12=2 7+10=17 2/17

OpenStudy (anonymous):

@killswitch1811 don't just give answers...

OpenStudy (anonymous):

x≤2/17

OpenStudy (anonymous):

@amber517 Almost... you messed up on the math just a bit. start back on the line: \[\large -10x+12x \le 7+10\]

OpenStudy (anonymous):

if i do -10+7it give me -3

OpenStudy (anonymous):

\[\large 12x - 10x \le 17\]\[\large 2x \le 17\]

OpenStudy (anonymous):

i'm not getting it there becuase -10+7=-3

OpenStudy (anonymous):

I think you've messed up on the math a bit. Do you wanna start from the beginning again?

OpenStudy (anonymous):

yes thats fine with me

OpenStudy (anonymous):

-10(x + 1) ≤ -12x + 7 -10x -10+10+12x ≤cancel -12x +7+10+12x -10x+12x≤7+10 -10x+12=2 7+10=17 x≤2/17

OpenStudy (anonymous):

we start off with: -10(x + 1) ≤ -12x + 7 Multiply -10 through: -10x -10 ≤ -12x +7 add 10 to both sides -10x ≤ -12x +7 +10 add 12x to both sides 12x - 10x ≤ 7+10

OpenStudy (anonymous):

12x-10x=2x

OpenStudy (anonymous):

7+10=17

OpenStudy (anonymous):

then you move the two over and that brings it to 17/2

OpenStudy (anonymous):

Right! then divide by two and you'll have your finished inequality. :D

OpenStudy (anonymous):

Good job!

OpenStudy (anonymous):

ok thanks